# A Survey on Large Language Models with Multilingualism: Recent Advances and New Frontiers

### v1

@@ 9.1 Bias Categories
The bias of LLMs in the multilingual scenario can be divided into two categories: language bias and demographic bias Xu et al. (2024a). Intuitively, the former is due to the imbalance of available training corpus for different languages Wei et al. (2023a); Nicholas and Bhatia (2023), where English possesses the most text corpus Dodge et al. (2021); Penedo et al. (2024); Gao et al. (2020); Soldaini et al. (2024), resulting in the performance degradation of the LLMs when generalized to other language settings Zhang et al. (2023j, 2024c). The demographic bias occurs due to embedded biases and misinformation on the internet, leading to LLMs unavoidably inheriting demographic biases across gender, race, and political backgrounds Ferrara (2023); Wang et al. (2022c); Levy et al. (2023); Yeh et al. (2023). This exacerbates existing inequalities, perpetuates stereotypes, and reinforces discrimination.
The core of multilingual LLM research is to improve the language modeling ability in other medium/low-resource languages while maintaining competence in English. Concerning demographic biases, previous attempts to mitigate bias and align LLMs with human values have primarily concentrated on English Hutchinson et al. (2020); Nadeem et al. (2021); Wan et al. (2023). Consequently, bias and ethical issues persist in other languages, potentially leading to significant negative impacts for non-English-speaking users.
@@
RTP-LX de Wynter et al. (2024) 28 Bias, Identity attack, Insult Microaggression, Self-harm Sexual content, Toxicity, Violence Small LM Large LM Exact Label Match Interrater Reliability de Wynter et al. (2024)
MGB Yu et al. (2024) zh,en,de,pt,es Gender Bias Masked LM MBE, MGL, LSG, MSG Yu et al. (2024)
MIBs Zhao et al. (2020) en,es,de,fr Gender Bias, Occupation Bias Word Embedding inBias Zhao et al. (2020)
MozArt Piqueras and Søgaard (2022) en,de,es,fr Gender Bias, Language Masked LM Close Test
@@ 9.2 Multilingual Debias
Language bias of LLMs persists in the multilingual scenario as a consequence of the dominance of English resources and the insufficiency of other languages on the internet. To enhance the model ability on low-resource languages, a common practice is to incorporate large-scale data Gao et al. (2020); Dodge et al. (2021) for training. The extensive training data facilitates language transfer, especially among typologically similar languages. Furthermore, the strategies such as curriculum learning Wei et al. (2023a) and up-sampling Lin et al. (2022); Shliazhko et al. (2022); Labrak et al. (2024) progressively increase the proportion of non-English resource. These techniques expose LLMs to a wider range of languages while maximizing the utilization of existing data.
To mitigate demographic bias in the multilingual scenario, Zhao et al. Zhao et al. (2020) extend word embedding bias to the cross-lingual and Piqueras et al. Piqueras and Søgaard (2022) evaluate group bias of three pre-trained LM (mBERT, XLM-R, and mT5) on four languages (En, Es, De, and Fr). Besides, Vashishtha et al. Vashishtha et al. (2023) extend debiasing strategies such as counterfactual data augmentation and self-debias to non-English languages, revealing a greater potential for debiasing and generalization among linguistically similar languages. However, they only investigate a few Indian languages, without comprehensive mitigation strategies for broader language groups. For a more comprehensive evaluation, De et al. de Wynter et al. (2024) introduce RTP-LX, a dataset designed for identifying culture-specific toxic languages with much wider coverage (28 languages and 8 different classes). Experimental results on up-to-date LLMs (Mistral Jiang et al. (2023), Gemma Team et al. (2024), GPT-4 Yuan et al. (2024), etc) demonstrate that even the popular models still struggle to judge history or content-dependent toxic content. Moreover, Lin et al. Lin et al. (2022) and Shliazhko et al. Shliazhko et al. (2022) analyze safety and bias in multilingual PLMs (XGLM and mGPT), respectively. They observe that multilingual PLMs pronounce gender bias in certain occupations, while few-shot learning has minimal impact on performance improvement. To analyze gender bias robustly, Yu et al. Yu et al. (2024) propose a novel model-based approach to generate sentence pairs. Based on mBERT Devlin et al. (2018), Reusens et al. Reusens et al. (2023) investigate cross-lingual transferability of debias techniques on 4 languages and stimulate cross-lingual debiasing effectiveness with additional pre-training.
@@ 9.3 Limitations and Discussion
To mitigate the cross-lingual bias is a promising research question and it has profound implications for the welfare, fairness, and esteem of numerous social and racial groups. Due to the scarcity of high-quality data in low-resource languages and the absence of pertinent evaluation benchmarks, effective bias detection and elimination remain largely unexplored, underscoring the necessity and imperative for future research. Besides, machine translation can mitigate the issue of low-resource languages via pseudo data generation Prabhumoye et al. (2018); Hoang et al. (2018) but it omits the cultural context or fails to capture the cultural nuances of a specific language Brislin (1970). The local and cultural backgrounds are critical to prejudice and hate speech. Thus, leveraging raw corpus in the local original context is important for detecting toxic content by native speakers.
@@ 10 Conclusion and Future Directions
In this paper, we summarize the existing representative research efforts on LLMs in the multilingual scenario from multiple perspectives. We first rethink the transitions between previous and current research on pre-trained language models. Based on the main aspect, the survey is divided into several sections from the view of training paradigms, inference strategies, security, multi-domain, data resources, and benchmark evaluation. Besides, we appeal to the research community for bias and fairness when exploring multilingual models. We also discuss several urgent challenges related to each investigated aspect and provide reflections and potential solutions for future work. Finally, considering the rapid growth of LLM research, we establish a continuously updated repository to provide relevant literature with the latest advancements of LLMs in the multilingual scenario.

### v2

@@ 10.1 Bias Categories
The bias of LLMs in the multilingual scenario can be divided into two categories: language bias and demographic bias [28]. Intuitively, the former is due to the imbalance of available training corpus for different languages [43, 457], where English possesses the most text corpus [462, 474, 475, 476], resulting in the performance degradation of the LLMs when generalized to other language settings [477, 478]. The demographic bias occurs due to embedded biases and misinformation on the internet, leading to LLMs unavoidably inheriting demographic biases across gender, race, and political backgrounds [479, 480, 481, 482]. This exacerbates existing inequalities, perpetuates stereotypes, and reinforces discrimination.
The core of multilingual LLM research is to improve the language modeling ability in other medium/low-resource languages while maintaining competence in English. Concerning demographic biases, previous attempts to mitigate bias and align LLMs with human values have primarily concentrated on English [483, 484, 485]. Consequently, bias and ethical issues persist in other languages, potentially leading to significant negative impacts for non-English-speaking users.
@@
RTP-LX [486] 28 Bias, Identity attack, Insult Microaggression, Self-harm Sexual content, Toxicity, Violence Small LM Large LM Exact Label Match Interrater Reliability [486]
MGB [487] zh,en,de,pt,es Gender Bias Masked LM MBE, MGL, LSG, MSG [487]
MIBs [488] en,es,de,fr Gender Bias, Occupation Bias Word Embedding inBias [488]
MozArt [489] en,de,es,fr Gender Bias, Language Masked LM Close Test
@@ 10.2 Multilingual Debias
Language bias of LLMs persists in the multilingual scenario as a consequence of the dominance of English resources and the insufficiency of other languages on the internet. To enhance the model ability on low-resource languages, a common practice is to incorporate large-scale data [475, 462] for training. The extensive training data facilitates language transfer, especially among typologically similar languages. Furthermore, the strategies such as curriculum learning [43] and up-sampling [38, 37, 156] progressively increase the proportion of non-English resource. These techniques expose LLMs to a wider range of languages while maximizing the utilization of existing data.
To mitigate demographic bias in the multilingual scenario, Zhao et al. [488] extend word embedding bias to the cross-lingual and Piqueras et al. [489] evaluate group bias of three pre-trained LM (mBERT, XLM-R, and mT5) on four languages (En, Es, De, and Fr). Besides, Vashishtha et al. [490] extend debiasing strategies such as counterfactual data augmentation and self-debias to non-English languages, revealing a greater potential for debiasing and generalization among linguistically similar languages. However, they only investigate a few Indian languages, without comprehensive mitigation strategies for broader language groups. For a more comprehensive evaluation, De et al. [486] introduce RTP-LX, a dataset designed for identifying culture-specific toxic languages with much wider coverage (28 languages and 8 different classes). Experimental results on up-to-date LLMs (Mistral [46], Gemma [491], GPT-4 [148], etc) demonstrate that even the popular models still struggle to judge history or content-dependent toxic content. Moreover, Lin et al. [38] and Shliazhko et al. [37] analyze safety and bias in multilingual PLMs (XGLM and mGPT), respectively. They observe that multilingual PLMs pronounce gender bias in certain occupations, while few-shot learning has minimal impact on performance improvement. To analyze gender bias robustly, Yu et al. [487] propose a novel model-based approach to generate sentence pairs. Based on mBERT [271], Reusens et al. [492] investigate cross-lingual transferability of debias techniques on 4 languages and stimulate cross-lingual debiasing effectiveness with additional pre-training.
@@ 10.3 Limitations and Discussion
To mitigate the cross-lingual bias is a promising research question and it has profound implications for the welfare, fairness, and esteem of numerous social and racial groups. Due to the scarcity of high-quality data in low-resource languages and the absence of pertinent evaluation benchmarks, effective bias detection and elimination remain largely unexplored, underscoring the necessity and imperative for future research. Besides, machine translation can mitigate the issue of low-resource languages via pseudo data generation [493, 494] but it omits the cultural context or fails to capture the cultural nuances of a specific language [495]. The local and cultural backgrounds are critical to prejudice and hate speech. Thus, leveraging raw corpus in the local original context is important for detecting toxic content by native speakers.
@@ 11 Conclusion and Future Directions
In this paper, we summarize the existing representative research efforts on LLMs in the multilingual scenario from multiple perspectives. We first rethink the transitions between previous and current research on pre-trained language models. Based on the main aspect, the survey is divided into several sections from the view of training paradigms, inference strategies, information retrieval, security, multi-domain, data resources, and benchmark evaluation. Besides, we appeal to the research community for bias and fairness when exploring multilingual models. We also discuss several urgent challenges related to each investigated aspect and provide reflections and potential solutions for future work. Finally, considering the rapid growth of LLM research, we establish a continuously updated repository to provide relevant literature with the latest advancements of LLMs in the multilingual scenario.